The following data are the monthly salaries y and the grade point averages x for students who obtained a bachelors degree in business administration. GPA 2.6 3.4 3.7 3.1 3.5 2.9 Monthly Salary (S) 3,600 3,800 4,200 3,700 4,100 2,500 The estimated regression equation for these data is y- 564.3 +964.3x and MSE-268,482. a. Develop a point estimate of the starting salary for a student with a GPA of 3.0 (to 1 decimal). b. Develop a 95% confidence interval for the mean starting salary for all students with a 3.0 GPA (to 2 decimals). C. Develop a 95% prediction interval for Ryan Dailey, a student with a GPA of 3.0 (to 2 decimals).
Answer #1

:

GPA (x) Monthly salary (y) GPA*Monthly salary GPA2 Monthly salary2
2.6 3600 9360 6.76 12960000
3.4 3800 12920 11.56 14440000
3.7 4200 15540 13.69 17640000
3.1 3700 11470 9.61 13690000
3.5 4100 14350 12.25 16810000
2.9 2500 7250 8.41 6250000
Sum = 19.2 21900 70890 62.28 81790000

ar x = Sigma x/n = 19.2/6 = extbf{3.2}

(Σ.r)2 = 62.28-19.22/6 = 0.84

Regression Equation:

у = 564.3 964.32.

MSE = 268482

a) Point estimate of the starting salary for a student with a GPA of 3.0

hat y^* = 564.3 + 964.3 *3 = $ extbf{ 3457.2}

b) At alpha = 0.05 and df = 6-2 = 4, t_{alpha/2} = 2.776

95% Confidence interval for the mean starting salary for all student with a GPA of 3.0

left (hat y^* pm t_{alpha/2}*sqrt{MSE* left (rac{1}{n} + rac{(x^*-ar x)^2}{SS_{XX}} ight )} ight )

left (3457.2 pm 2.776*sqrt{(268482)*left ( rac{1}{6} + rac{(3-3.2)^2}{0.84} ight )} ight )

left (3457.2 pm 2.776*239.858 ight )

left ( extbf{ 2791.35, 4123.05} ight )

c) 95% Prediction interval for Ryan dailey, a student with a GPA of 3.0

left (hat y^* pm t_{alpha/2}*sqrt{MSE* left (1+rac{1}{n} + rac{(x^*-ar x)^2}{SS_{XX}} ight )} ight )

left (3457.2 pm 2.776*sqrt{(268482)*left ( 1+rac{1}{6} + rac{(3-3.2)^2}{0.84} ight )} ight )

left (3457.2 pm 2.776*570.9762 ight )

left ( extbf{ 1872.17, 5042.23} ight )